OpenStudy (anonymous):

Solve x divided by negative six is less than or equal to three for x. Show your work.

OpenStudy (campbell_st):

so you have \[\frac{x}{-6} \le 3\] multiply both sides of the inequality by -6 and reverse the inequality for the solution

OpenStudy (anonymous):

anyone know?

OpenStudy (anonymous):

you multiply -6 by itself and 3 by -6\[\frac{ x }{ -6\times-6 }=3\times-6\]

OpenStudy (anonymous):

so it would be \[x \le -18\]

OpenStudy (anonymous):

@mercedesbustos12

OpenStudy (campbell_st):

well not quite its \[\frac{x}{-6} \times -6 \le 3 \times -6\] then reverse the inequality

OpenStudy (campbell_st):

and don't forget to reverse the inequality this is why using @minimissy22 \[-24 \le -18\] but is \[\frac{-24}{-6} \le 3\]

OpenStudy (anonymous):

so which one is it?

OpenStudy (anonymous):

@campbell_st @minimissy22

OpenStudy (anonymous):

It is @campbell_st I realized I messed up sorry @mercedesbustos12